The expression that is equivalent to( 5x3+5 )2 /25x3

1
Expert's answer
2021-04-28T05:09:38-0400

(5x3+5)2/(25x3) = 52*(x8)2/(25x3 )= 25x16/(25x3 )= x16/x3 = x13